OpenStudy (anonymous):

Calculus Help! PLease! Find an equation of the tangent line to the graph of f at the given point F(x)=1/1+x, (0,1)

OpenStudy (anonymous):

find \[f \prime (x)\] then \[f \prime (0)\] then substitute in \[y-y _{1}=f \prime(x)(x-x _{1})\]

OpenStudy (anonymous):

using \[f/g=(f*g \prime-g*f \prime)/g^2\]

OpenStudy (anonymous):

@lilly21 start the derivative by apply quotient rule :)

OpenStudy (anonymous):

I did okay this is how i started....f(x+h)-f(x)/h..........1/x+1-1/x+h all over h then..i did this 1-1(1)/x+h+1

OpenStudy (anonymous):

but there is a rule for that,on top